OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{n^{2}+1}{n^{3} +1}\] my book did a limit comparison test for this \[\lim_{n\rightarrow\infty} \frac{a_{n}}{b_{n}}\] and somehow they took an n/1 and multiplied it to \[\frac{n^{2}+1}{n^{3} +1}\] and made this sum divergent, i believe its supposed to be convergent

OpenStudy (experimentx):

\[ \frac{n^{2}+1}{n^{3} +1} = \frac{n^{2}}{n^{3} +1} + \frac{1}{n^{3} +1} \approx \frac{1}{n} + \frac{1}{n^{3} +1}\] 1/n diverges, and 1/n^3+1 converges.

OpenStudy (anonymous):

how does \[\frac{n^{2}}{n^{3} +1} \approx \frac{1}{n} \]

OpenStudy (anonymous):

i guess dividing numerator and denominator by 1/n^2

OpenStudy (experimentx):

as n tends to infinity, \[ \lim_{n \rightarrow \infty } \frac{1}{n + \frac 1 {n^2}}\] Yup ..

OpenStudy (experimentx):

if you want you can perform integral test .. i don't think ratio test will be conclusive. also try pluggin in wolf

OpenStudy (anonymous):

ok so 1/n is divergent ...why? \[1/\infty\] should be convergent

OpenStudy (experimentx):

why do you think ... ?? ratio test in inconclusive ... from integral test you find 1/n is divergent. You can prove it shortly if you want. like -1+1-1+1-1+1 ..

OpenStudy (anonymous):

ok i'll use the integral test

OpenStudy (experimentx):

Let 1/n converge to x say 1+1/2+1/3+ 1/4 + 1/5 +1/6 + ... = x 1/2 + 1/4 + 1/6 + .... 1 + 1/3 +1/5 + 1/7 + ... = x 1/2(x) + 1 + 1/3 +1/5 + 1/7 + ... = x or, 1 + 1/3 +1/5 + 1/7 + ... = 1/2 (x) = 1/2 + 1/4 + 1/6 + ... which is a contradiction ... so 1/n is divergent.

OpenStudy (anonymous):

i see...back to the limit comparison test for a second...what about this \[\lim_{n\rightarrow\infty} \frac{a_{n}}{b_{n}}\] ordeal

OpenStudy (anonymous):

where \[a_{n} =\frac{n^{2}}{n^{3} +1}\] and \[b_{n}= 1/n\]

OpenStudy (anonymous):

sorry an =\[\frac{n^{2}+1}{n^{3} +1} \]

OpenStudy (experimentx):

 take n tends to infinity ... try pluggin in wolframalpha ... but i don't think ratio test will be conclusive.